OpenStudy (robellio123):

How can I answer this???? PLEASE HELP Julie divided x^4 + 3x^3 -9x^2 +6 by a polynomial. Her answer was x^3 - 2x^2 + x - 5 with a remainder of 31. What polynomial did Julie divide by?

OpenStudy (daniyil_the_spy):

ok

OpenStudy (mww):

We have P(x) =A(x)Q(x) + R(x) and want to find A(x) P(x)/A(x) = Q(x) + R(x)/A(x) = (A(x)Q(x) + R(x))/A(x) expand the numerator i.e. A(x)Q(x)+ R(x) but first let A(x) = (x - a) for some constant a and compare coefficients against P(x)

OpenStudy (3mar):

 I think this is how Julie has done this!